OpenStudy (ashy98):

Simplify completely quantity x squared minus 10 x minus 24 all over x squared minus 3 x minus 108 and find the restrictions on the variable.

OpenStudy (ashy98):

A.) \[\frac{ x+2 }{ x+9} ,x \neq-2,x \neq9\] B.) \[\frac{ x+2 }{x+9 } ,x \neq-9,x \neq12\] C.) \[\frac{ x-2 }{ x-9 },x \neq2,x \neq-9\] D.) \[\frac{ x-2 }{ x-9 },x \neq-9,x \neq12\]

OpenStudy (ashy98):

I think the answer is A

OpenStudy (mathstudent55):

If you know how to use Latex, then why do you spell out the expression in the post? Why not write, for example, \(x^2 + 5x - 3\), which is much faster to read?

OpenStudy (ashy98):

its b, right not a